# Jaime Fernández Fisac and Chen Lv

### v2

@@ Abstract
Existing traffic simulation models often fail to capture the complexities of real-world scenarios, limiting the effective evaluation of autonomous driving systems. We introduce Versatile Behavior Diffusion (VBD), a novel traffic scenario generation framework that utilizes diffusion generative models to predict scene-consistent and controllable multi-agent interactions in closed-loop settings. VBD achieves state-of-the-art performance on the Waymo Sim Agents Benchmark and can effectively produce realistic and coherent traffic behaviors with complex agent interactions under diverse environmental conditions. Furthermore, VBD offers inference-time scenario editing through multi-step refinement guided by behavior priors and model-based optimization objectives. This capability allows for controllable multi-agent behavior generation, accommodating a wide range of user requirements across various traffic simulation applications. Despite being trained solely on publicly available datasets representing typical traffic conditions, we introduce conflict-prior and game-theoretic guidance approaches that enable the creation of interactive, long-tail safety-critical scenarios—essential for comprehensive testing and validation of autonomous vehicles. Lastly, we provide in-depth insights into effective training and inference strategies for diffusion-based traffic scenario generation models, highlighting best practices and common pitfalls. Our work significantly advances the ability to simulate complex traffic environments, offering a powerful tool for the development and assessment of autonomous driving technologies. Project website: https://sites.google.com/view/versatile-behavior-diffusion.
keywords:
Autonomous Driving, Traffic Simulation, Multi-agent Interaction, Scenario Generation, Diffusion Model
1 1. Introduction
@@
Comprehensive testing is crucial for ensuring the safe deployment of autonomous driving systems. Traditionally, these assessments rely on replaying logged trajectories from driving datasets in simulation environments (Li et al., 2022). However, because other traffic participants in the replay do not react to the autonomous vehicle’s actions, log-replay methods often fail to account for interactive traffic scenarios, such as pedestrians stepping into crosswalks or vehicles negotiating right-of-way at four-way stops. This limitation introduces a significant simulation-to-reality gap, where simulated test results may not translate to real-world performance, undermining the validity of simulation testing.
@@
Figure 1: Illustration of the VBD framework. The model consists of a query-centric Transformer scene condition encoder, a marginal multi-modal behavior predictor for individual agents, and a diffusion generator for joint multi-agent behavior. In training, the model only utilizes commonly used driving datasets collected from public roads. In testing, the model’s versatility is showcased through various purposes enabled by simple sampling and different guidance structures (e.g., prior guidance for sophisticated target agent’s behavior control and game-theoretic cost guidance for adversarial behaviors). The VBD model functions as a generalized simulation engine and operates effectively in a closed-loop traffic simulator.
@@
To address this challenge, traffic simulators that incorporate reactive traffic behavior models, which account for interactions between agents and their surroundings, have become increasingly popular. These models enable the synthesis of scenarios that better reflect the complex dynamics of real-world traffic, including lane changes, merges, and responses to traffic controls. Conventional heuristic-based traffic modeling approaches (Treiber et al., 2000; Kesting et al., 2007) often fail to scale to the complexity of realistic traffic maneuvers, such as subtle navigation around construction zones, pedestrians stepping into crosswalks, or vehicles inching forward at congested intersections. As a result, recent efforts (Shi et al., 2022; Igl et al., 2022; Nayakanti et al., 2023; Zhang et al., 2023c; Feng et al., 2023) have increasingly shifted towards data-driven methods that leverage large-scale driving datasets (Caesar et al., 2020; Wilson et al., 2023; Montali et al., 2023) and behavior cloning (BC) techniques to more accurately model traffic behaviors. Despite these advancements, existing methods still focus on predicting marginal open-loop trajectories for individual agents, potentially resulting in a lack of scene consistency. This means that when multiple agents’ most likely trajectories are rolled out together, unnatural interactions and even collisions can occur (Chen et al., 2022), compromising the realism and reliability of simulated scenarios. One common approach to improve scene consistency is by controlling agents’ motions in a receding horizon fashion or utilizing an auto-regressive policy (Zhou et al., 2024; Wu et al., 2024), relying on the model to re-plan and adapt to each other’s actions. However, by controlling all agents with a single shared policy, the planned trajectories often lack diversity. Consequently, generated scenarios struggle to effectively model complex, joint interactions among multiple agents. Additionally, this practice may be inefficient, as the computational requirements scale linearly with the number of traffic participants in the scene. Therefore, scenario generation models in traffic simulators should be designed for scene-centric prediction, where future trajectories of all agents are jointly generated within a unified and interactive framework, enabling more scalable and realistic simulation of complex traffic scenarios.
@@
Furthermore, traffic simulators should offer versatility and enable controllable scenario generation to satisfy both explicit (e.g., goals or targets) and implicit (e.g., collision avoidance) requirements. Existing models are often specialized in generating nominal behaviors conditioned on high-level objectives, such as navigation goals (Zhang et al., 2023c). However, scenario generation models should also address the “long-tail” challenge in most datasets, where rare and dangerous situations are underrepresented (Liu and Feng, 2024), potentially leading to inadequate testing in critical scenarios. Previous works in safety-critical traffic scenario generation (Cao et al., 2022; Hanselmann et al., 2022; Rempe et al., 2022; Zhang et al., 2023a) have tailored training objectives to encourage adversarial behaviors (e.g., inducing collisions with the ego agent), which limits their ability to capture the nuances of real-world traffic interactions. For instance, a model trained solely to generate aggressive driving behaviors may struggle to produce scenarios where surrounding agents attempt to avoid such adversarial actions, highlighting the constraints imposed by its learning objectives.
Therefore, a more generalized traffic simulation framework is needed to produce a wide spectrum of scenarios accommodating various requirements at inference time without necessitating retraining or — a level of flexibility that current models frequently lack. This flexibility is crucial in traffic simulation, as it allows users to assign different modes to agents, enables the evaluation of edge cases, and facilitates the exploration of novel situations, ultimately providing a more comprehensive and customizable simulation experience. To achieve both realism and controllability in multi-agent behavior modeling, one potential solution is to first learn an optimization objective that captures nominal driving behavior, and then align scenarios to user specifications by solving an optimization problem that combines the learned objectives with additional heuristics. This approach involves modeling optimization objectives — a standard problem in inverse optimal control (IOC) or inverse reinforcement learning (IRL) — which aims to infer the underlying cost function or reward structure governing an agent’s behavior (Ng et al., 2000; Ziebart et al., 2008). Existing IOC/IRL methods have been applied to modeling trajectories of traffic agents, ranging from directly mapping scenarios to costs (Rosbach et al., 2019; Wang et al., 2023a; Huang et al., 2021) to learning cost weights for a set of handcrafted heuristics through differentiable optimization layers (Huang et al., 2023c, a; Diehl et al., 2023). However, these methods have not matched the performance achieved by behavior cloning methods (Shi et al., 2022; Nayakanti et al., 2023; Wu et al., 2024) in capturing realistic traffic behaviors.
@@
To overcome these limitations, we leverage diffusion models (also known as score-based models) (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Song et al., 2020c; Ho et al., 2020), a class of generative modeling methods that gradually recover structured data from random noise. We develop Versatile Behavior Diffusion (VBD), a flexible and scalable model for traffic behavior generation. To the best of our knowledge, VBD is the first open-source framework capable of generating realistic, interactive, and controllable traffic scenarios for any number of agents in a closed-loop setting. As illustrated in Figure 1, despite being trained only on commonly used public road traffic datasets, VBD can accommodate a variety of tasks—such as normal or adversarial behavior generation, counterfactual reasoning, and reactive simulation—through guided sampling combined with user-specified objectives. Our model demonstrates state-of-the-art performance on the large-scale simulated agent benchmark, as shown in Figure 2, and offers versatile generation capabilities. Additionally, we elucidate both training and inference-time design choices of VBD to provide valuable insights for future work. The primary contributions of this paper are threefold:
@@
    We propose Versatile Behavior Diffusion for generalized traffic simulation, capable of generating realistic and controllable traffic agent behaviors with excellent closed-loop simulation performance.
@@
    We conduct extensive empirical studies to investigate the effects of various training and inference settings on multi-agent behavior generation using diffusion models, providing valuable insights for future research.
In the rest of this manuscript, we first establish a conceptual connection between diffusion generative modeling and imitation learning for multi-agent traffic behavior modeling, as well as compositional scenario generation in Section 3. Then, we detail the model architecture and training methodology of VBD in Section 4. Section 5 illustrates that VBD can generate realistic and interactive traffic scenarios without guidance in closed-loop simulations. Furthermore, we demonstrate multiple real-world use cases of VBD to enhance generation quality and sample user-desired behaviors through guidance in Section 6, and generate safety-critical scenarios using guidance in Section 7. Finally, we ablate important design choices necessary for high-quality generation in Section 8.
2 2. Related Work
2.1 2.1 Traffic Simulation
@@
Traditional model-based or heuristic-based methods (Treiber et al., 2000; Kesting et al., 2007) often fail to capture complex agent interactions in real-world scenarios, resulting in a significant gap between simulation and reality. Consequently, there has been a growing shift towards learning-based methods to enhance the realism and interactivity of traffic simulations (Suo et al., 2021). For example, BITS (Xu et al., 2023b) employs imitation learning to simulate agent behaviors by inferring high-level intentions and replicating low-level driving actions. The socially-controllable behavior generation model proposed in Chang et al. (2023b) focuses on simulating social and interactive behaviors. Symphony (Igl et al., 2022) integrates learning-based policies with parallel beam search to further enhance realism. TrafficBots (Zhang et al., 2023c) introduces a shared policy conditioned on specific goals to generate configurable behaviors. Trajeglish (Philion et al., 2023) introduces a multi-agent sequence of motion tokens using a GPT-like encoder-decoder architecture, achieving state-of-the-art realism metrics. Another research direction focuses on generating safety-critical or adversarial scenarios to test the robustness of driving systems. STRIVE (Rempe et al., 2022) generates challenging scenarios that can induce collisions with the ego planner through optimization in latent space. Similarly, AdvDO (Cao et al., 2022) and KING (Hanselmann et al., 2022) utilize optimization-based methods to generate adversarial trajectories for robust planning. CAT (Zhang et al., 2023a) selects conflicting trajectories from agents’ predicted behavior distributions to construct adversarial scenarios, supporting robust training of RL driving agents. However, most existing simulation models lack versatility, as they are specifically trained either for normal behaviors (maximum likelihood) or adversarial scenarios. This limitation restricts their applicability in general-purpose traffic simulators. We aim to bridge this gap by developing a generative framework that supports various tasks and requirements to enable generalizable, controllable, and realistic traffic simulation.
2.2 2.2 Behavior Prediction
@@
Behavior or trajectory prediction models are widely utilized in both traffic simulations and driving policies (Zhang et al., 2023c; Sun et al., 2023; Feng et al., 2023). Recent advances in learning-based behavior prediction models have significantly improved the accuracy of predicting single-agent motions (Huang et al., 2022; Nayakanti et al., 2023; Zhou et al., 2023) as well as joint multi-agent interactions at the scene level (Mo et al., 2022; Shi et al., 2023; Huang et al., 2023b). Leveraging a large amount of real-world data, these models are capable of generating multi-modal distributions of possible behaviors for multiple agents in a scene. However, these prediction models often face limitations in scalability and efficiency when applied to large-scale scenarios with numerous agents. Therefore, more powerful generative models, including diffusion models (Jiang et al., 2023; Niedoba et al., 2023) and autoregressive Transformer models (Seff et al., 2023; Zhou et al., 2024), have been increasingly applied in behavior prediction and generation tasks, demonstrating superior performance in large-scale predictions and simulations. Our proposed VBD model integrates multi-modal behavior prediction as high-level intention priors, enabling the estimation of realistic behavior distributions and sophisticated control over target agents’ behaviors, which facilitates the diffusion model in generating specific scenarios based on the intended behaviors of particular agents.
2.3 2.3 Diffusion Models for Traffic Simulation
@@
Diffusion models, also known as score-based models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song et al., 2020c, a), have gained widespread popularity in a variety of generative tasks, including image (Zhang et al., 2023b), audio (Kong et al., 2020), and video (Esser et al., 2023) generation. Recently, their application has extended to traffic scenario generation, where their ability to generate diverse and controllable outputs is particularly valuable. For instance, SceneDM (Guo et al., 2023) utilizes a diffusion model to generate consistent joint motions for all agents in a scene. MotionDiffuser (Jiang et al., 2023) employs a diffusion-based representation for joint multi-agent motion prediction and introduces a constrained sampling framework for controlled trajectory sampling. CTG (Zhong et al., 2023b) combines diffusion modeling with signal temporal logic (STL) rules to ensure compliance with traffic rules in generated trajectories. CTG++ (Zhong et al., 2023a) leverages Large Language Models (LLMs) to translate user queries into loss functions, guiding the diffusion model toward generating query-compliant scenarios. TRACE (Rempe et al., 2023) proposes a guided diffusion model to generate future trajectories for pedestrians, employing analytical loss functions to impose trajectory constraints. Other models, such as DiffScene (Xu et al., 2023a) and Safe-Sim (Chang et al., 2023a), adopt guided diffusion techniques with adversarial optimization to generate safety-critical traffic scenarios. Concurrent work from Waymo (Jiang et al., 2024) also explored the idea of generating joint agent trajectories for traffic scenario generation. They proposed to use amortized diffusion to improve the quality and computational efficiency for closed-loop roll-outs. While our work is closely related to these guided diffusion models, these approaches often remain limited to specific tasks and lack versatility across different simulation applications. Additionally, there has been limited conceptual exploration of diffusion models within the context of traffic simulation. We aim to advance this area by proposing training strategies for diffusion-based behavior generation models and exploring various inference strategies to enhance both realism and versatility.
3 3. Problem Formulation
3.1 3.1 Traffic Scenario Generation as Optimization
Consider a traffic scenario $\mathcal{S}=({\mathbf{x}},\mathbf{u},\mathbf{c})$ with an episode length ${T}$ containing a tensor of ${A}$ agents’ trajectories ${\mathbf{x}}=({\mathbf{x}}^{1},\dots,{\mathbf{x}}^{A})\in\mathbb{R}^{{A}\times{T}\times{D_{x}}}$ and control sequences $\mathbf{u}=(\mathbf{u}^{1},\cdots,\mathbf{u}^{A})\in\mathbb{R}^{{A}\times{T}\times{D_{u}}}$. The context of the scene $\mathbf{c}\in\mathbb{R}^{D_{c}}$ includes information regarding the road map, traffic light status, and the initial joint state of all agents ${x}_{0}$, etc. Given an optimization objective $\mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c})$, we formulate scenario generation as a finite-horizon optimal control problem by:
@@
$\begin{split}&\underset{\mathbf{u}\in\mathbb{R}^{{A}\times{T}\times{D_{u}}}}{\min}\leavevmode\nobreak\ \mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c}),\\ \text{s.t.}\quad&{\mathbf{x}}_{0}={x}_{0},\\ &{\mathbf{x}}_{{t}+1}={f}({\mathbf{x}}_{{t}},\mathbf{u}_{{t}}),\leavevmode\nobreak\ \forall{t}\in\{0,\ldots,{T}-1\},\end{split}$ (1)
where ${f}$ represents the discrete-time joint dynamics. If our goal is to generate realistic (statistically representative) scenarios, the objective $\mathcal{J}_{\theta}$ should incentivize real-world driving behaviors, which can be achieved through statistical loss metrics or inverse reinforcement learning (IRL). Once $\mathcal{J}_{\theta}$ is established, we need to tractably find an optimal joint control sequence $\mathbf{u}$ through numerical optimization or reinforcement learning.
3.2 3.2 Generative Modeling as Trajectory Optimization
Instead of solving the aforementioned IL problem in two steps, prior works (Ho and Ermon, 2016; Finn et al., 2016) have shown a strong connection between IL and generative modeling under a generative-adversarial training framework. Extending the analysis by Du and Mordatch (2019), Liu et al. (2022), and Chi et al. (2023), we show that synthesizing a diffusion generative model in this IL setting can be viewed as learning the gradient descent step of a particular optimal control solver.
Consider a dataset $\mathbb{D}$ with scenario triplets sampled independently from an unknown distribution $p$. Since we are interested in scenario generation given a scene context and the recorded trajectory ${\mathbf{x}}$ as a control sequence $\mathbf{u}$ under known dynamics ${f}$, we can factorize the probability density function as $p(\mathcal{S})=p(\mathbf{u}|\mathbf{c})p(\mathbf{c})$. Under the Maximum Entropy IRL (Ziebart et al., 2008) formulation, we aim to approximate $p(\mathbf{u}|\mathbf{c})$ as the Boltzmann distribution of an optimization objective:
@@
where $Z_{\theta}$ is the partition function. Equation 2 resembles the Energy-Based Models (EBM) (LeCun et al., 2006; Song and Kingma, 2021). Specifically, we want to learn the parameter $\theta$ of the optimization objective that maximizes the conditional log-likelihood of the dataset $\mathbb{D}$:
$\theta=\operatorname*{arg\,max}_{\hat{\theta}}\ {\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta}}(\mathbf{u}|\mathbf{c})].$ (3)
Ideally, we can employ score-matching (Hyvärinen and Dayan, 2005; Vincent, 2011; Song and Ermon, 2019; Song et al., 2020b) to directly learn the gradient of $\mathcal{J}_{\theta}$ w.r.t the control (our random variable of interest) as the score function:
$\nabla_{\mathbf{u}}\log p(\mathbf{u}|\mathbf{c})$ $\approx\mathbf{s}_{\theta}(\mathbf{u}|\mathbf{c})$
$:=\nabla_{\mathbf{u}}\log p_{\theta}(\mathbf{u}|\mathbf{c})$ (4)
[…]
If $\nabla_{\mathbf{u}}\mathcal{J}_{\theta}$ was obtained over the entire action space, we could use it for gradient descent. However, since the dataset contains mostly near-optimal scenarios, the gradient estimation in suboptimal regions of the action space (away from demonstration data) may be inaccurate or not well-defined. To overcome this issue, a class of approaches (Du and Mordatch, 2019; Song et al., 2020b; Ho et al., 2020; Song et al., 2020c, a) utilize a stochastic process to gradually diffuse $p$ into noised distributions $p_{k}$ for $k$ steps until it becomes a known distribution $p_{K}=\pi$. These methods are commonly known as Diffusion models (Ho et al., 2020; Song et al., 2020a) and are later generalized as score-based models by Song et al. (2020c). Specifically, we train a step-conditioned score function $\mathbf{s}_{\theta}(\tilde{\mathbf{u}}|\mathbf{c},k)$ to approximate the gradient of the log noised distribution $\nabla_{\tilde{\mathbf{u}}}\log p_{k}(\tilde{\mathbf{u}})$ by:
$\theta=\operatorname*{arg\,max}_{\hat{\theta}}$ ${\mathbb{E}}_{\mathcal{S}\sim\mathbb{D},k\sim{\mathcal{U}}(0,K)}{\mathbb{E}}_{\tilde{\mathbf{u}}\sim p_{k}(\cdot|\mathbf{u})}\big{[}$ (5)
$\lambda(k)\lVert\nabla_{\tilde{\mathbf{u}}}\log p_{k}(\tilde{\mathbf{u}}|\mathbf{u})-\mathbf{s}_{\hat{\theta}}(\tilde{\mathbf{u}}|\mathbf{c},k)\rVert\big{]},$
where $\lambda(k)$ is a positive weighting function. At inference time, we can generate scenarios by first randomly selecting $\tilde{\mathbf{u}}$ from the known distribution $\pi$ and sampling through the reverse diffusion process.
Connecting this formulation of generative modeling with trajectory optimization, we can view the forward diffusion as uplifting original data distribution into a higher-dimensional space augmented by diffusion step $k$. By injecting noise, we achieve good coverage over the entire action space in the final step $K$ so that $\mathbf{s}_{\theta}(\tilde{\mathbf{u}}|\mathbf{c},K)$ are well defined for random $\tilde{\mathbf{u}}$. Sampling through reverse diffusion can be interpreted as stochastic gradient descent towards high-probability regions with a fixed descent direction along the diffusion step, analogous to the direct shooting method in optimal control. We note that at low noise level $k$, as $p_{k}$ is close to the original data distribution $p$, $\mathbf{s}_{\theta}(\mathbf{u}|\mathbf{c},k\rightarrow 0)\approx-\nabla_{\mathbf{u}}\mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c})$, which is the gradient we originally try to model. Therefore, the generative modeling of scenarios can be viewed as an explicit solution of IL by learning the gradient steps of trajectory optimization and solving the optimal control problem through reverse diffusion sampling.
In the remainder of the paper, without loss of generality, we consider a specific form of Diffusion model, diffusion-denoising probabilistic models (DDPM) (Ho et al., 2020), which is also known as the discrete-time variance-preserving score-based SDE (VP-SDE) (Song et al., 2020c). The equivalence between the original DDPM training objective and the score-matching loss (Equation 5) has been shown in Luo (2022).
3.3 3.3 Denoising Diffusion Probabilistic Model
The discrete-time formulation of the forward diffusion process of DDPM (Nichol and Dhariwal, 2021) can be described as:

### v3

# Versatile Behavior Diffusion for Generalized Traffic Agent Simulation

@@ Abstract
Existing traffic simulation models often fall short in capturing the intricacies of real-world scenarios, particularly the interactive behaviors among multiple traffic participants, thereby limiting their utility in the evaluation and validation of autonomous driving systems. We introduce Versatile Behavior Diffusion (VBD), a novel traffic scenario generation framework based on diffusion generative models that synthesizes scene-consistent, realistic, and controllable multi-agent interactions. VBD achieves strong performance in closed-loop traffic simulation, generating scene-consistent agent behaviors that reflect complex agent interactions. A key capability of VBD is inference-time scenario editing through multi-step refinement, guided by behavior priors and model-based optimization objectives, enabling flexible and controllable behavior generation. Despite being trained on real-world traffic datasets with only normal conditions, we introduce conflict-prior and game-theoretic guidance approaches. These approaches enable the generation of interactive, customizable, or long-tail safety-critical scenarios, which are essential for comprehensive testing and validation of autonomous driving systems. Extensive experiments validate the effectiveness and versatility of VBD and highlight its promise as a foundational tool for advancing traffic simulation and autonomous vehicle development. Project website: https://sites.google.com/view/versatile-behavior-diffusion.
I Introduction
@@
Simulation testing is crucial for ensuring the safe deployment of autonomous driving systems. Traditionally, these assessments rely on replaying logged trajectories from driving datasets in simulation environments [32]. However, because other traffic participants in the replay do not react to the autonomous vehicle’s actions, log-replay methods often fail to account for interactive traffic scenarios, such as pedestrians stepping into crosswalks or vehicles negotiating right-of-way at four-way stops. This limitation introduces a significant simulation-to-reality gap, where simulated test results may not translate to real-world performance, undermining the credibility of simulation-based testing.
@@
Figure 1: Illustration of the VBD framework. The model consists of a query-centric Transformer scene condition encoder, a marginal multi-modal behavior predictor for individual agents, and a diffusion generator for joint multi-agent behavior. In training, the model only utilizes commonly used traffic datasets. In testing, the model’s versatility is showcased through various purposes enabled by simple sampling and different guidance structures (e.g., prior guidance for sophisticated target agents’ behavior control and game-theoretic cost guidance for adversarial behaviors). VBD functions as a generalized simulation engine and operates effectively in a closed-loop traffic simulator.
@@
To address this challenge, traffic simulators that incorporate reactive traffic behavior models have become increasingly popular. These models account for interactions between agents and their surroundings, enabling the synthesis of scenarios that better reflect the complex dynamics of real-world traffic. However, conventional heuristic-based traffic modeling approaches [54, 29] often fail to scale to the complexity of realistic traffic maneuvers. As a result, recent efforts [45, 25, 37, 62, 15] have shifted towards data-driven methods. These methods leverage large-scale driving datasets [2, 56, 36] and behavior cloning to model traffic behaviors. Despite these advancements, existing methods still focus on predicting marginal open-loop trajectories for individual agents, potentially resulting in a lack of scene consistency. This means that when multiple agents’ most likely trajectories are rolled out together, unnatural interactions and even collisions can occur [7], compromising the realism and reliability of simulated scenarios.
@@
Furthermore, traffic simulators should offer versatility and enable controllable scenario generation to satisfy both explicit (e.g., goals or targets) and implicit (e.g., collision avoidance) requirements. Existing models are often specialized in generating nominal behaviors conditioned on high-level objectives, such as navigation goals [62]. However, scenario generation models should also address the long-tail challenge in most datasets, where rare and dangerous situations are underrepresented [33], potentially leading to inadequate testing in critical scenarios. Previous works in safety-critical traffic scenario generation [3, 19, 42, 60] have employed tailored training objectives to encourage adversarial behaviors (e.g., inducing collisions with the ego agent), which limits their ability to capture the nuances of real-world traffic interactions. For instance, a model trained to generate aggressive driving behaviors may struggle to produce scenarios where surrounding agents attempt to avoid such adversarial actions, highlighting the constraints imposed by its specialized learning objectives. Therefore, a more generalized traffic simulation framework is needed to produce a wide spectrum of scenarios accommodating various requirements at inference time without retraining. This flexibility is crucial in traffic simulation, as it allows users to assign different behavioral modes to agents, enables the evaluation of edge cases, and facilitates the exploration of long-tail situations, providing a more comprehensive and customizable simulation tool.
@@
To overcome these limitations, we leverage diffusion models [47, 49, 51, 20], a class of generative modeling methods that gradually recover structured data from random noise. We develop Versatile Behavior Diffusion (VBD), a flexible and scalable model for traffic behavior generation. VBD is capable of generating realistic, interactive, and controllable traffic scenarios for a large number of agents. As illustrated in Fig. 1, despite being trained only on commonly used public road traffic datasets, VBD can accommodate a variety of tasks, such as normal or adversarial behavior generation, counterfactual reasoning, and reactive simulation, through guided sampling combined with user-specified objectives. Our VBD model demonstrates strong performance on the Waymo Sim Agent benchmark and offers versatile generation capabilities. Our main contributions are summarized as follows:
@@
    We propose Versatile Behavior Diffusion for generalized traffic simulation, capable of generating realistic and controllable traffic agent behaviors with strong closed-loop simulation performance.
@@
    We conduct ablation studies to investigate the effects of various training and inference settings on multi-agent behavior generation.
II Related Work
II-A Traffic Simulation
@@
Traditional model-based or heuristic-based methods [54, 29] often fail to capture complex agent interactions in real-world scenarios, resulting in a significant gap between simulation and reality. Consequently, there has been a growing shift towards learning-based methods to enhance the realism and interactivity of traffic simulations [53]. For example, BITS [59] employs imitation learning to simulate agent behaviors by inferring high-level intentions and replicating low-level driving actions. The socially-controllable behavior generation model proposed in [6] focuses on simulating social and interactive behaviors. Symphony [25] integrates learning-based policies with parallel beam search to further enhance realism. TrafficBots [62] introduces a shared policy conditioned on specific goals to generate configurable behaviors. Trajeglish [40] introduces a multi-agent sequence of motion tokens using a GPT-like encoder-decoder architecture. Another research direction focuses on generating safety-critical or adversarial scenarios to test the robustness of driving systems. STRIVE [42] generates challenging scenarios that can induce collisions with the ego planner through optimization in latent space. Similarly, AdvDO [3] and KING [19] utilize optimization-based methods to generate adversarial trajectories for robust planning. CAT [60] selects conflicting trajectories from agents’ predicted behavior distributions to construct adversarial scenarios, supporting robust training of RL driving agents. However, most existing simulation models lack versatility, as they are specifically trained either for normal behaviors (maximum likelihood) or adversarial scenarios. We aim to bridge this gap by developing a generative framework that supports various tasks and requirements to enable generalizable, controllable, and realistic traffic simulation.
II-B Behavior Prediction
@@
Behavior or trajectory prediction models are widely utilized in both traffic simulations and driving policies [62, 52, 15]. Recent advances in learning-based behavior prediction models have significantly improved the accuracy of predicting single-agent motions [24, 37, 67] as well as joint multi-agent interactions at the scene level [35, 46, 23]. Leveraging large amounts of real-world data, these models are capable of generating multi-modal distributions of possible behaviors for multiple agents in a scene. However, these prediction models often face limitations in scalability and efficiency when applied to large-scale scenarios with numerous agents. Therefore, more powerful generative models, including diffusion models [27, 39] and autoregressive Transformer models [44, 66], have been increasingly applied in behavior prediction and generation tasks, demonstrating superior performance in large-scale predictions and simulations. Our proposed VBD model integrates multi-modal behavior prediction as high-level intention priors, enabling the estimation of realistic behavior distributions and sophisticated control over target agents’ behaviors. This integration facilitates the diffusion model in generating specific scenarios based on the intended behaviors of particular agents.
II-C Diffusion Models for Traffic Simulation
@@
Diffusion models [49, 20, 51, 48] have gained widespread popularity in a variety of generative tasks, including image [61], audio [30], and video [13] generation. Recently, their application has been extended to traffic scenario generation. For instance, SceneDM [18] utilizes a diffusion model to generate consistent joint motions for all agents in a scene. MotionDiffuser [27] employs a diffusion-based representation for joint multi-agent motion prediction and introduces a constrained sampling framework for controlled trajectory sampling. CTG [65] combines diffusion modeling with signal temporal logic (STL) rules to ensure compliance with traffic rules in generated trajectories. CTG++ [64] leverages Large Language Models (LLMs) to translate user queries into loss functions, guiding the diffusion model toward generating query-compliant scenarios. TRACE [41] proposes a guided diffusion model to generate future trajectories for pedestrians, employing analytical loss functions to impose trajectory constraints. Other models, such as DiffScene [58] and Safe-Sim [5], adopt guided diffusion techniques with adversarial optimization to generate safety-critical traffic scenarios. SceneDiffuser [28] proposed to use amortized diffusion to improve the quality and computational efficiency for closed-loop rollouts. While our work is closely related to these guided diffusion models, these approaches often remain limited to specific tasks and lack versatility across different simulation applications.
III Preliminaries
III-A Problem Formulation
We formulate traffic scenario generation as a conditional trajectory optimization problem in a multi-agent setting. A scenario $\mathcal{S}=({\mathbf{x}},\mathbf{u},\mathbf{c})$ consists of agent trajectories ${\mathbf{x}}=({\mathbf{x}}^{1},\dots,{\mathbf{x}}^{A})\in\mathbb{R}^{{A}\times{T}\times{D_{x}}}$, corresponding control sequences $\mathbf{u}=(\mathbf{u}^{1},\dots,\mathbf{u}^{A})\in\mathbb{R}^{{A}\times{T}\times{D_{u}}}$, and scene context $\mathbf{c}\in\mathbb{R}^{D_{c}}$ that encodes maps, traffic signals, and initial agent states ${x}_{0}$.
We can define the scenario generation task as a finite-horizon optimal control problem:
@@ III-A Problem Formulation
$\begin{split}\underset{\mathbf{u}}{\min}~&\mathcal{J}_{\theta}({\mathbf{x}},\mathbf{u};\mathbf{c}),\\ \text{s.t.}\quad&{\mathbf{x}}_{0}={x}_{0},\\ &{\mathbf{x}}_{{t}+1}={f}({\mathbf{x}}_{t},\mathbf{u}_{t}),\ \forall{t}\in{0,\dots,{T}-1}\end{split}$ (1)
where f{f} defines discrete-time joint dynamics, and $\mathcal{J}_{\theta}$ promotes realistic, feasible behaviors.
To model the distribution over control sequences $\mathbf{u}$ conditioned on context $\mathbf{c}$, we adopt the Maximum Entropy IRL [68] formulation, representing the target distribution as a Boltzmann distribution:
@@
which resembles an Energy-Based Model (EBM) [31, 50]. The learning objective is to maximize the conditional likelihood of $\mathbf{u}$ under the model $p_{\theta}$:
$\theta=\operatorname*{arg\,max}_{\hat{\theta}}~{\mathbb{E}}_{\mathcal{S}\sim\mathbb{D}}[\log p_{\hat{\theta}}(\mathbf{u}|\mathbf{c})].$ (3)
Directly estimating $\nabla_{\mathbf{u}}\log p_{\theta}(\mathbf{u}|\mathbf{c})$ (the score function) is challenging due to limited coverage in low-likelihood regions. To address this, recent works have employed diffusion models [20, 51, 48], which gradually transform data samples into noise through a forward stochastic process and then learn to reverse the process. Training such a model can be interpreted as learning the gradient steps of trajectory optimization in the score space, thus bridging generative modeling and control. As shown by [12, 34, 9], reverse diffusion sampling in this context approximates stochastic gradient descent toward low-cost, high-likelihood trajectories. This insight allows diffusion models to serve as both generative models and implicit planners that solve the control problem defined in Eq. 1.
III-B Denoising Diffusion Probabilistic Model (DDPM)
The discrete-time formulation of the forward diffusion process of DDPM [38] can be described as: